OpenStudy (tootsi123):

Which of the following is polynomial factor of 8x^2+22x-6? A (4x-3) B (4x+3) C (8x-2) D (8x+3) E (8x+6)

OpenStudy (tootsi123):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

what is the GCF of 8,22, and 6 ?

OpenStudy (tootsi123):

Oh wait i think i get it hold on... 8= 2^3 22=2*11 6=2*3

OpenStudy (tootsi123):

is that right so far??

jimthompson5910 (jim_thompson5910):

notice how 2 is a factor in each, so we can factor out 2 8x^2+22x-6 = 2(4x^2+11x - 3)

jimthompson5910 (jim_thompson5910):

how would we factor 4x^2+11x - 3 ?

OpenStudy (tootsi123):

Wouldn't you start simplifying the problem?

jimthompson5910 (jim_thompson5910):

multiply the first coefficient (4) and the last term (-3) 4*(-3) = -12 what two factors of -12 add to 11 ?

jimthompson5910 (jim_thompson5910):

put another way: what two numbers multiply to -12 and add to 11?

OpenStudy (tootsi123):

Wait i am confused... i thought you couldn't combine terms unless they were like terms.... and would 3 be one

jimthompson5910 (jim_thompson5910):

I'm not combining like terms

jimthompson5910 (jim_thompson5910):

I'm asking for the two numbers because it will help factor

OpenStudy (tootsi123):

oh. Okay so 11 is a prime number so its only factor is one.

OpenStudy (tootsi123):

-12 could be 6,2,3,4

jimthompson5910 (jim_thompson5910):

here are all of the ways to multiply two whole numbers to get -12 -1*12 -2*6 -3*4 -4*3 -6*2 -12*1 which pair of factors will add up to 11?

OpenStudy (tootsi123):

The first one

jimthompson5910 (jim_thompson5910):

yep -1*12 = -12 -1 + 12 = 11

jimthompson5910 (jim_thompson5910):

so we can break up the 11x into -1x + 12x

OpenStudy (tootsi123):

Oh okay...

jimthompson5910 (jim_thompson5910):

4x^2+11x - 3 4x^2-1x + 12x - 3

jimthompson5910 (jim_thompson5910):

then we'd factor 4x^2-1x + 12x - 3 by grouping

jimthompson5910 (jim_thompson5910):

group up the terms in pairs 4x^2-1x + 12x - 3 (4x^2-1x) + (12x - 3) do you see what step is next?

OpenStudy (tootsi123):

Simplify them.

jimthompson5910 (jim_thompson5910):

specifically you factor each group

jimthompson5910 (jim_thompson5910):

how can you factor 4x^2-1x ?

OpenStudy (tootsi123):

I don't know?

jimthompson5910 (jim_thompson5910):

we can factor out x, agreed? 4x^2 - 1x = x(4x - 1)

OpenStudy (tootsi123):

oh yeah since they are like terms

jimthompson5910 (jim_thompson5910):

how would you factor 12x - 3 ?

OpenStudy (tootsi123):

Wouldn't you just do 12 -3 and then add the x

jimthompson5910 (jim_thompson5910):

what factors do 12 and 3 have in common?

jimthompson5910 (jim_thompson5910):

12x and -3 aren't like terms

OpenStudy (tootsi123):

oh duh 3

jimthompson5910 (jim_thompson5910):

yeah so 12x - 3 = 3(4x - 1)

jimthompson5910 (jim_thompson5910):

4x^2-1x + 12x - 3 (4x^2-1x) + (12x - 3) x(4x - 1) + 3(4x - 1)

jimthompson5910 (jim_thompson5910):

what factor shows up twice?

OpenStudy (tootsi123):

one

jimthompson5910 (jim_thompson5910):

what else

OpenStudy (tootsi123):

4

jimthompson5910 (jim_thompson5910):

the 4x-1 shows up twice

jimthompson5910 (jim_thompson5910):

we can factor that out x(4x - 1) + 3(4x - 1) will turn into (x+3)(4x - 1)

jimthompson5910 (jim_thompson5910):

so hopefully you see how 8x^2+22x-6 factors to 2(x+3)(4x - 1)

OpenStudy (tootsi123):

I think so

jimthompson5910 (jim_thompson5910):

go over the steps a few times to make sure you have a solid grasp of it. you can ask questions if you aren't sure about a step or two

OpenStudy (tootsi123):

Okay